# {cutout}02.753cm0cm1 empty ViC-Bench     : Benchmarking Visual-Interleaved Chain-of-Thought Capability in MLLMs with Free-Style Intermediate State Representations

### v2

@@ 5 Experiments
GPT-4o-20240513 [47] OpenAI $\times$ 26.40 2.00 28.60 100 2.00 31.42 0.85 86.90
o3 [48] OpenAI 34.80 18.00 53.27 98.80 24.00 58.22 1.38 93.93
o4-mini [49] OpenAI 32.40 0.80 16.00 88.40 1.60 12.49 -0.32 35.98
Gemini-2.5-pro-preview-03-25 [63] Google 38.40 20.80 57.00 100 28.40 64.09 0.13 97.17
Gemini-2.0-flash [62] Google 11.60 5.60 41.67 99.93 10.80 39.17 -0.53 72.13
Claude-3.5-Sonnet-20241022 [3] Anthropic 34.40 0.80 26.60 99.87 0.80 9.34 -1.55 23.89
Step-1o-vision-32k [58] StepFun $\times$ 32.20 0.00 20.87 98.87 1.20 22.93 -0.01 92.70
Doubao-1.5-vision-pro-32k [7] ByteDance $\times$ 22.80 0.00 16.73 100 0.00 16.10 0.07 88.97
SenseChat-Vision-V5.5 [54] SenseTime $\times$ 26.40 1.60 32.87 98.80 1.60 31.91 -0.26 97.26
Hunyuan-Vision-20250103 [68] Tencent $\times$ 24.80 0.00 16.80 98.80 0.00 12.10 -1.30 69.63
@@
LLaVA-OneVision-72B [32] ByteDance & NTU $\times$ 14.40 0.00 17.27 100 0.00 15.80 0.00 100
InternVL-2.5-78B [12] Shanghai AI Lab $\times$ 27.20 2.00 31.73 100 2.40 28.25 -0.27 89.24
InternVL-3-78B [87] Shanghai AI Lab $\times$ 24.80 0.80 26.60 100 4.40 33.24 -0.20 84.85
Qwen2-VL-72B-Instruct [70] Alibaba $\times$ 25.20 0.40 19.47 100 0.40 19.48 0.06 99.61
Qwen2.5-VL-72B-Instruct [5] Alibaba $\times$ 34.00 0.40 25.13 99.93 0.80 25.75 -0.33 94.25
Qwen-Omni-Turbo [67] Alibaba $\times$ 20.80 0.00 17.33 99.60 0.00 16.88 -0.14 90.62
GLM-4V-Plus [64] Zhipu AI $\times$ 34.80 0.40 22.07 100 0.80 14.36 0.00 61.74
SkyWork-R1V1-38B [50] Skywork 20.40 0.00 15.13 80.80 0.00 14.60 0.00 74.89
5.1 Experimental Setup
Evaluation MLLMs. We select 18 top-performing MLLMs for comprehensive evaluation. Regarding proprietary models, we include powerful OpenAI and Gemini series, such as o3/o4 [49] and Gemini-2.5-pro [63]. We also test popular models from the competitive organizations, such as Claude-3.5-Sonnet [3] and Doubao [7]. As for open-source models, we select the respresentive series, such as InternVL and Qwen-VL [87, 5, 64]. Besides, we select the varying versions of a model family with its largest capacity to investigate performance gaps. Finally, we evaluate MLLMs with targeted thinking capabilities, such as SkyWork-R1V1-38B [50]. Note that we take GPT-4o [47] as the baseline model.
Implementation Details. We access proprietary and open models via APIs and local deployment. The maximum token limit is 8192, temperature is 0, both Top-K and Top-P are 1. The rest of hyper-parameter settings keeps same with VLMEvalKit [20]. All the experiments are conducted on a machine with 8 $\times$ NVIDIA A100 GPUs (80G). Due to space constraint, we only apply IPII strategy for maze navigation here. Specifically, Level-1 prompts are consistent with origin prompts, which are illustrated in Fig. 7. Level-2 builds upon Level-1 by incorporating the following prompting information: Please make sure that after executing the move at each step, you should update your internal intermediate visual state, rather than remaining in the initial input visual state, as displayed in Fig. 11. As shown in Fig. 12, Level-3 further builds upon Level-2 by adding the coordinates information of the starting and end points in the maze navigation task.
Table 4: The performace evaluations of advanced MLLMs on embodied long-horizon planning.
Method Organization Reason- oriented Stage-1 Stage-2
ACC ACC Recall_o
@@ 5.1 Experimental Setup
GPT-4o-20240513 [47] OpenAI $\times$ 55.60 0.80 2.68
o3 [48] OpenAI 62.80 1.60 2.52
o4-mini [49] OpenAI 62.00 0.40 2.25
Gemini-2.5-pro-preview-03-25 [63] Google 60.80 18.80 25.21
Gemini-2.0-flash [62] Google 42.80 4.80 9.06
Claude-3.5-Sonnet-20241022 [3] Anthropic 54.80 0.40 1.48
Step-1o-vision-32k [58] StepFun $\times$ 66.80 4.00 5.64
Doubao-1.5-vision-pro-32k [7] ByteDance $\times$ 41.60 1.60 1.80
SenseChat-Vision-V5.5 [54] SenseTime $\times$ 30.80 0.80 1.72
Hunyuan-Vision-20250103 [68] Tencent $\times$ 30.40 2.80 7.96
@@
LLaVA-OneVision-72B [32] ByteDance & NTU $\times$ 50.00 2.40 3.54
InternVL-2.5-78B [12] Shanghai AI Lab $\times$ 11.60 0.40 0.47
InternVL-3-78B [87] Shanghai AI Lab $\times$ 29.20 2.00 2.98
Qwen2-VL-72B-Instruct [70] Alibaba $\times$ 53.20 1.60 2.99
Qwen2.5-VL-72B-Instruct [5] Alibaba $\times$ 61.20 1.20 2.61
Qwen-Omni-Turbo [67] Alibaba $\times$ 40.40 0.80 1.29
GLM-4V-Plus [64] Zhipu AI $\times$ 47.20 5.60 9.05
SkyWork-R1V1-38B [50] Skywork 27.20 0.40 1.26
Table 5: The performace evaluations of advanced MLLMs on complex counting.
@@
GPT-4o-20240513 [47] OpenAI $\times$ 28.80 12.00 44.30 14.40 41.50 7.60
o3 [48] OpenAI 38.40 10.40 40.00 17.60 50.70 10.10
o4-mini [49] OpenAI 40.00 2.40 25.40 3.20 24.70 6.20
Gemini-2.5-pro-preview-03-25 [63] Google 41.20 18.80 36.60 18.00 43.60 0.00
Gemini-2.0-flash [62] Google 46.80 5.60 22.00 12.80 33.50 1.40
Claude-3.5-Sonnet-20241022 [3] Anthropic 21.60 3.20 22.00 4.40 28.90 6.80
Step-1o-vision-32k [58] StepFun $\times$ 43.20 9.60 28.50 6.80 27.80 0.00
Doubao-1.5-vision-pro-32k [7] ByteDance $\times$ 43.60 13.20 33.20 6.00 21.80 -11.00
SenseChat-Vision-V5.5 [54] SenseTime $\times$ 24.00 14.80 44.70 15.20 44.10 5.10
Hunyuan-Vision-20250103 [68] Tencent $\times$ 32.00 8.00 30.90 3.20 16.00 -0.30
@@
LLaVA-OneVision-72B [32] ByteDance & NTU $\times$ 35.20 6.00 22.60 8.00 24.20 -1.70
InternVL-2.5-78B [12] Shanghai AI Lab $\times$ 21.20 6.40 23.00 4.80 22.70 0.30
InternVL-3-78B [87] Shanghai AI Lab $\times$ 28.80 11.60 34.90 15.20 37.00 0.00
Qwen2-VL-72B-Instruct [70] Alibaba $\times$ 36.00 11.60 36.20 12.40 33.00 -0.60
Qwen2.5-VL-72B-Instruct [5] Alibaba $\times$ 50.00 8.00 32.90 10.00 29.70 0.00
Qwen-Omni-Turbo [67] Alibaba $\times$ 27.60 16.40 44.00 10.00 26.60 -7.90
GLM-4V-Plus [64] Zhipu AI $\times$ 37.20 14.40 43.00 9.20 37.10 0.00
SkyWork-R1V1-38B [50] Skywork 36.80 6.40 27.30 7.60 26.70 0.00
5.2 Main Results
@@
Maze Navigation. Tab. 2 displays that most MLLMs exhibit competent performance in Stage 1. Performance significantly drops in Stage 2, indicating that current MLLMs have limitations in open-ended spatial reasoning and perception. In Stage 3, with the supports of free-style VIS, all models consistently achieves gains in global-level ACC and fine-grained $\mathbf{R\_o}$, leading to impressive ThinkGain, which indicates the effectiveness of free-style IVS in tackling deficiencies of spatial-aware cognition. However, we observe decline in Legality, which indicates that external knowledge could further confuse the thinking trajectories of weaker models. From Fig. 4, we find that response length is inversely proportional to efficiency in maze navigation. These outcomes suggest that free-style IVS can enhance performance in spatial reasoning for MLLMs with strong learned reasoning priors.
@@
Jigsaw Puzzle. Tab. 3 indicates that there also exists significant declines from Stage 1 to 2, confirming that open-ended QA poses challenges to visual CoT in this task, which is likely to stem from the difficulties in understanding AIGC semantics and distinguishing semantically incoherent patches. MLLMs in Stage 3 promoted by free-style IVS exhibit global-level gains compared to Stage 2, but surprisingly achive negative ThinkGain with impressive drops in Legality, which could be due to the irrational IVS utiliztions leading to invalid decisions. From Fig. 4, we figure out that top-performing models [48, 63] have concise responses, while SkyWork-R1V1 [50] with longer responses exhibits wipeout. These failures shows the deficiencies of current models in using free-form IVS, reflecting the inadequate VI-CoT. Future works should explore promoting positive utilizations of IVS in training.
@@
Figure 4: Quantitative comparisons on averge response length, ACC, and ThinkGain for three tasks.
Figure 5: The further investigations with IPII strategy in VI-CoT capability for maze navigation.
Embodied Long-Horizon Planning. Most models exhibit significant declines from Stage 1 to 2, particularly Step-1o [58]. We attribute this observation to models exploiting multiple-choice QA by matching semantics with the given sequence to make hacking decision. However, results in Stage 2 demonstrate that models do not truly understand the instructions and engage in reasoning. Instead, they tend to directly make plans based on semblances. For Stage 3, we select the real-world scenario for qualitative comparison, as shown in Fig. 6. We conclude that powerful models can grasp implicit instructions to perform reasoning for object searching, indicating the benefits of free-style IVS in facilitating embodied planning.
@@
Complex Counting. ACC in Stage 2 is inferior to Stage 1, which confirms our conclusions in above tasks. In Stage 3, some models [7, 67, 68] exhibit ACC declines and negative ThinkGain. Surprisingly, Gemini-2.5-pro [63] also exhibits drops with null ThinkGain. These outcomes indicate that MLLMs still suffer from object hallucinations and deficiencies in basic perception, leading to accumulated erroneous trajectories. In case studies, we also find that IVS assist models in recognizing heads, yet they still make wrong decisions, consistent with observations in LLMs [22, 78]. From Fig. 4, we find that models with relatively better performance tend to have shorter responses. The above results reveal the impressive difficulties MLLMs face with counting, which should be promoted by developing targeted representations or training strategies.
Further Investigations with IPII Strategy. Proprietary models in Stage 1 consistently achieve better performance. Among open models, [87, 5] experience varying degrees of decline. Surprisingly, SkyWork-R1V1 [50] finally reaches 98.40% ACC, indicating that R1-like MLLMs might have higher sensitivities to prompts. In Stage 2, all models do not achieve impressive outcomes across levels, revealing that existing MLLMs might lack the capability to implicitly update their internal IVS through textual CoT in the open-ended QA. These results also indicates the necessity of utilizing free-style IVS as the external knowledge to improve reasoning. In Stage 3, performance is closely related to the inherent VI-CoT priors. Strong models [63] can leverage external knowledge to achieve better results, while weaker models might be confused by extra knowledge, leading to drops.

### v3

@@ V-A Experimental Setup
GPT-4o-20240513 [openai2024gpt4o] OpenAI $\times$ 26.40 2.00 28.60 100 2.00 31.42 0.85 86.90
o3 [openai2025o3mini] OpenAI 34.80 18.00 53.27 98.80 24.00 58.22 1.38 93.93
o4-mini [openai2025o4mini] OpenAI 32.40 0.80 16.00 88.40 1.60 12.49 -0.32 35.98
Gemini-2.5-pro-preview-03-25 [Gemini-2.5-pro-preview-03-25] Google 38.40 20.80 57.00 100 28.40 64.09 0.13 97.17
Gemini-2.0-flash [Gemini-2.0-flash] Google 11.60 5.60 41.67 99.93 10.80 39.17 -0.53 72.13
Claude-3.5-Sonnet-20241022 [anthropic2024claude3d5] Anthropic 34.40 0.80 26.60 99.87 0.80 9.34 -1.55 23.89
Step-1o-vision-32k [step-1o-vision-32k] StepFun $\times$ 32.20 0.00 20.87 98.87 1.20 22.93 -0.01 92.70
Doubao-1.5-vision-pro-32k [Doubao-1.5-vision-pro-32k] ByteDance $\times$ 22.80 0.00 16.73 100 0.00 16.10 0.07 88.97
SenseChat-Vision-V5.5 [SenseChat-Vision-V5.5] SenseTime $\times$ 26.40 1.60 32.87 98.80 1.60 31.91 -0.26 97.26
Hunyuan-Vision-20250103 [Hunyuan-Vision-20250103] Tencent $\times$ 24.80 0.00 16.80 98.80 0.00 12.10 -1.30 69.63
@@
LLaVA-OneVision-72B [Llava-onevision] ByteDance & NTU $\times$ 14.40 0.00 17.27 100 0.00 15.80 0.00 100
InternVL-2.5-78B [Internvl] Shanghai AI Lab $\times$ 27.20 2.00 31.73 100 2.40 28.25 -0.27 89.24
InternVL-3-78B [InternVL3] Shanghai AI Lab $\times$ 24.80 0.80 26.60 100 4.40 33.24 -0.20 84.85
Qwen2-VL-72B-Instruct [Qwen2-vl] Alibaba $\times$ 25.20 0.40 19.47 100 0.40 19.48 0.06 99.61
Qwen2.5-VL-72B-Instruct [Qwen2d5-vl] Alibaba $\times$ 34.00 0.40 25.13 99.93 0.80 25.75 -0.33 94.25
Qwen-Omni-Turbo [Qwen-Omni-Turbo] Alibaba $\times$ 20.80 0.00 17.33 99.60 0.00 16.88 -0.14 90.62
GLM-4V-Plus [glm2024chatglm] Zhipu AI $\times$ 34.80 0.40 22.07 100 0.80 14.36 0.00 61.74
SkyWork-R1V1-38B [Skywork-r1v] Skywork 20.40 0.00 15.13 80.80 0.00 14.60 0.00 74.89
TABLE IV: The performace evaluations of advanced MLLMs on embodied long-horizon planning.
Method Organization Reason- oriented Stage-1 Stage-2 Stage-3
ACC ACC Recall_o Legality ACC Recall_o ThinkGain Legality
@@
GPT-4o-20240513 [openai2024gpt4o] OpenAI $\times$ 57.20 0.04 0.40 71.13 12.80 35.07 22.95 43.74
o3 [openai2025o3mini] OpenAI 63.60 1.20 1.58 70.89 16.00 29.48 20.69 56.57
o4-mini [openai2025o4mini] OpenAI 68.80 2.40 3.07 70.77 22.80 41.61 26.14 53.76
Gemini-2.5-pro-preview-03-25 [Gemini-2.5-pro-preview-03-25] Google 67.60 6.80 16.69 65.00 22.80 34.84 19.43 53.84
Gemini-2.0-flash [Gemini-2.0-flash] Google 41.20 1.20 3.79 59.37 7.60 18.74 14.97 47.69
Claude-3.5-Sonnet-20241022 [anthropic2024claude3d5] Anthropic 49.20 0.00 0.11 42.88 15.00 43.75 27.08 59.51
Step-1o-vision-32k [step-1o-vision-32k] StepFun $\times$ 61.20 3.20 3.44 68.86 6.80 23.79 10.77 42.68
Doubao-1.5-vision-pro-32k [Doubao-1.5-vision-pro-32k] ByteDance $\times$ 64.00 0.00 50.37 70.97 14.00 31.17 19.91 49.54
SenseChat-Vision-V5.5 [SenseChat-Vision-V5.5] SenseTime $\times$ 25.60 0.00 0.40 64.91 0.00 0.18 0.08 0.26
Hunyuan-Vision-20250103 [Hunyuan-Vision-20250103] Tencent $\times$ 22.80 0.00 0.24 36.76 3.60 21.49 14.47 35.40
@@
LLaVA-OneVision-72B [Llava-onevision] ByteDance & NTU $\times$ 48.80 0.00 17.31 72.71 0.00 4.17 2.50 11.55
InternVL-2.5-78B [Internvl] Shanghai AI Lab $\times$ 39.20 0.00 0.76 53.83 0.00 3.69 2.94 94.72
InternVL-3-78B [InternVL3] Shanghai AI Lab $\times$ 47.20 0.80 1.58 43.64 0.80 10.40 7.50 82.98
Qwen2-VL-72B-Instruct [Qwen2-vl] Alibaba $\times$ 46.40 1.20 1.76 74.17 10.80 25.15 11.45 28.32
Qwen2.5-VL-72B-Instruct [Qwen2d5-vl] Alibaba $\times$ 52.00 0.80 1.12 69.74 10.00 23.75 9.64 35.84
Qwen-Omni-Turbo [Qwen-Omni-Turbo] Alibaba $\times$ 26.40 0.90 0.00 0.00 0.71 6.00 3.51 7.71
GLM-4V-Plus [glm2024chatglm] Zhipu AI $\times$ 46.80 2.40 4.88 54.50 4.80 25.51 13.37 29.58
SkyWork-R1V1-38B [Skywork-r1v] Skywork 23.20 0.40 0.76 7.52 0.00 0.27 0.13 0.34
TABLE V: The performace evaluations of advanced MLLMs on complex counting.
@@
GPT-4o-20240513 [openai2024gpt4o] OpenAI $\times$ 28.80 12.00 44.30 14.40 41.50 7.60
o3 [openai2025o3mini] OpenAI 38.40 10.40 40.00 17.60 50.70 10.10
o4-mini [openai2025o4mini] OpenAI 40.00 2.40 25.40 3.20 24.70 6.20
Gemini-2.5-pro-preview-03-25 [Gemini-2.5-pro-preview-03-25] Google 41.20 18.80 36.60 18.00 43.60 0.00
Gemini-2.0-flash [Gemini-2.0-flash] Google 46.80 5.60 22.00 12.80 33.50 1.40
Claude-3.5-Sonnet-20241022 [anthropic2024claude3d5] Anthropic 21.60 3.20 22.00 4.40 28.90 6.80
Step-1o-vision-32k [step-1o-vision-32k] StepFun $\times$ 43.20 9.60 28.50 6.80 27.80 0.00
Doubao-1.5-vision-pro-32k [Doubao-1.5-vision-pro-32k] ByteDance $\times$ 43.60 13.20 33.20 6.00 21.80 -11.00
SenseChat-Vision-V5.5 [SenseChat-Vision-V5.5] SenseTime $\times$ 24.00 14.80 44.70 15.20 44.10 5.10
Hunyuan-Vision-20250103 [Hunyuan-Vision-20250103] Tencent $\times$ 32.00 8.00 30.90 3.20 16.00 -0.30
@@
LLaVA-OneVision-72B [Llava-onevision] ByteDance & NTU $\times$ 35.20 6.00 22.60 8.00 24.20 -1.70
InternVL-2.5-78B [Internvl] Shanghai AI Lab $\times$ 21.20 6.40 23.00 4.80 22.70 0.30
InternVL-3-78B [InternVL3] Shanghai AI Lab $\times$ 28.80 11.60 34.90 15.20 37.00 0.00
Qwen2-VL-72B-Instruct [Qwen2-vl] Alibaba $\times$ 36.00 11.60 36.20 12.40 33.00 -0.60
Qwen2.5-VL-72B-Instruct [Qwen2d5-vl] Alibaba $\times$ 50.00 8.00 32.90 10.00 29.70 0.00
Qwen-Omni-Turbo [Qwen-Omni-Turbo] Alibaba $\times$ 27.60 16.40 44.00 10.00 26.60 -7.90
GLM-4V-Plus [glm2024chatglm] Zhipu AI $\times$ 37.20 14.40 43.00 9.20 37.10 0.00
SkyWork-R1V1-38B [Skywork-r1v] Skywork 36.80 6.40 27.30 7.60 26.70 0.00
V-B Main Results
@@
Maze Navigation. Tab. II indicates that most MLLMs exhibit competent performance in Stage 1. Performance significantly drops in Stage 2, indicating that current MLLMs have limitations in open-ended spatial reasoning and perception. In Stage 3, with the supports of free-style VIS, all models consistently achieves gains in global-level ACC and fine-grained $\mathbf{R\_o}$, leading to impressive ThinkGain, which indicates the effectiveness of free-style IVS in tackling deficiencies of spatial-aware cognition. However, we observe decline in Legality, which indicates that external knowledge could further confuse the thinking trajectories of weaker models. From Fig. 4, we can clearly observe that response length is inversely proportional to efficiency in maze navigation, implying that excessive verbosity often signals uncertainty or error accumulation. These outcomes suggest that free-style IVS can act as a critical visual anchor, allowing MLLMs with strong priors to verify intermediate states and significantly enhance the spatial-aware reasoning performance.
@@
Jigsaw Puzzle. Tab. III indicates that there also exists significant declines from Stage 1 to 2, confirming that open-ended QA poses challenges to visual CoT in this task, which is likely to stem from the difficulties in understanding AIGC semantics and distinguishing semantically incoherent patches. MLLMs in Stage 3 promoted by free-style IVS exhibit global-level gains compared to Stage 2, but surprisingly achieve negative ThinkGain with impressive drops in Legality, which could be due to the irrational IVS utilizations leading to invalid decisions. Fig. 4 indicates that top-performing models [openai2025o3mini, Gemini-2.5-pro-preview-03-25] achieve high efficiency with concise responses. In stark contrast, SkyWork-R1V1 [Skywork-r1v] suffers from performance wipeout despite its lengthy reasoning chains, indicating that verbose CoT does not guarantee success without effective grounding. These failures underscore the deficiencies of current MLLMs in effectively integrating free-form IVS, revealing an immature VI-CoT capability where visual feedback may distract rather than guide. Consequently, future works should explore training paradigms that can effectively align visual generation with reasoning goals to ensure the positive utilization of IVS.
@@
Figure 4: The visualized comparisons on averge response length, ACC, and ThinkGain for three tasks.
Embodied Long-Horizon Planning. Most models exhibit significant declines from Stage 1 to Stage 2, exemplified by GPT-4o plummeting from $57.2\%$ to $0.04\%$. This exposes that current MLLMs possess a sophisticated linguistic shell but lack grounded physical world capability, leading to embodied hallucinations that defy basic laws. Deprived of options, models degenerate into blind guessing, with Skywork-R1V1 [Skywork-r1v] even failing basic instruction constraints with $7.52\%$ Legality metric. Crucially, this performance collapse is generally reversed in Stage 3, particularly for advanced models. The substantial ThinkGain metric demonstrates that free-style IVS serves as a vital visual anchor, empowering capable models to transform blind hallucinations into verifiable actions, thereby reactivating the embodied reasoning potential dormant in text-only contexts. For Stage 3, we further conduct extensive qualitative analysis in a real-world scenario, as illustrated in Fig. 6 below. This task poses a significant challenge as the targeted object is occluded, requiring models to infer implicit sub-goals for object searching. As observed, proprietary MLLMs [openai2024gpt4o, Gemini-2.5-pro-preview-03-25] demonstrate robust reasoning-for-planning capabilities. They successfully decompose the high-level instruction into actionable steps, and exhibit self-correction behaviors when initial attempts fail. In contrast, open-source model [Qwen2d5-vl] struggles with long-horizon dependencies, which exhibits failure modes including: (1) invalid navigation planning, where the model attempts to navigate directly to an invisible target; (2) task deviation, such as interacting with irrelevant objects; and (3) recursive behavior, getting trapped in repetitive loops of opening and closing the microwave. These comparisons underscore the critical role of free-style IVS in bridging the gap between high-level instructions and embodied execution, enabling models to maintain logical consistency over long horizons, while also revealing the significant performance gaps between open-source and closed-source models.
@@
Complex Counting. ACC in Stage 2 is inferior to Stage 1, which confirms our conclusions in the above tasks and reinforces that the removal of option-based hints exposes the inability of models to perform autonomous enumeration. In Stage 3, some models [Doubao-1.5-vision-pro-32k, Qwen-Omni-Turbo, Hunyuan-Vision-20250103] exhibit ACC declines and negative ThinkGain. Surprisingly, Gemini-2.5-pro [Gemini-2.5-pro-preview-03-25] also exhibits drops with null ThinkGain. These outcomes indicate that MLLMs still suffer from object hallucinations and deficiencies in basic perception, where low-quality generated visuals act as noise rather than valid references, leading to accumulated erroneous trajectories. In case studies, we also find that IVS assist models in recognizing heads, yet they still make wrong decisions, consistent with observations in LLMs [counting1, counting2]. As can be seen from Fig. 4, we observe that models with relatively better performance tend to produce shorter, more concise responses. This trend suggests that capable models can efficiently identify and count objects without resorting to verbose, redundant reasoning often indicative of uncertainty. These outcomes highlight the significant challenges MLLMs still face with complex counting, particularly in maintaining spatial consistency during long-horizon enumeration. As a result, we claim that future research should focus on developing targeted representations, such as explicit visual markers, or specialized training strategies to enhance object discrimination and spatial awareness.
Figure 5: The further investigations with IPII strategy in VI-CoT capability of advanced MLLMs for maze navigation.
V-C Further Investigations with IPII Strategy.
Our analysis across the three stages of IPII strategy demonstrates the distinct capabilities and limitations of MLLMs. In Stage 1, proprietary models consistently exhibit superior performance stability, while open-source models [InternVL3, Qwen2d5-vl] show a decline with increasing difficulty. The notable exception of SkyWork-R1V1 [Skywork-r1v], which shows a sharp accuracy surge at Level 3, suggests that R1-like MLLMs might have higher sensitivities to prompts. In Stage 2, the widespread failure to achieve significant performance gains across all models highlights a critical bottleneck, i.e., existing MLLMs might lack the capability to implicitly update their internal IVS through textual CoT in the open-ended QA. These results also indicate the necessity of utilizing free-style IVS as external knowledge to improve reasoning. Finally, in Stage 3, the benefit of introducing external free-style IVS leads to a clear performance divergence. Advanced MLLMs [Gemini-2.5-pro-preview-03-25] can successfully leverage the visual contexts to boost planning accuracy, whereas weaker MLLMs often struggle with information overload, leading to confusion and decreased performance.